Exercise 2-15 (LO2-4) Mollys Candle Shop has several retail stores in the coastal areas of North and South Carolina. Many of Mollys customers ask her to ship their purchases. The following chart shows the number of packages shipped per day for the last 100 days. 28 30 20 10F5 23 18 13 10 5 10 15 20 25 30 35 Number of Packages a. What is this chart called?

What is the total number of frequencies?

What is the class interval?

What is the class frequency of the 25 up to 30 class?

What is the midpoint of the 20 up to 25 class?

On how many days were there 15 or more packages shipped?
Answer #1

a. This chart is called Histogram.

b. The total number of frequencies is 5+ 13 + 28 + 23 + 18 + 10 +3 =100

c. Class interval = 5

d. the class frequency of the 25 up to 30 class is 10

e. the midpoint of the 20 up to 25 class is (20 + 25)/2 = 22.5

f. 23 + 18 + 10 + 3 = 54

54 days were there 15 or more packages shipped